Any one can help me how to solve this problem . A moving van charge a flat rate of $25 per day plus .12 per mile for every mile over 100 driven . If Millie's bill was $29.46 how many miles to the nearest mile did she drive in all thanks
Answer by macston(5194)

Charged mileage=M=mileage-100 mi. (If mileage<100 mi, M=0)
Total Charge= $25/day+$0.12(M)=$29.46 Subtract $25 from each side
$25-$25+$0.12(M)=$29.46-$25
$0.12(M)=$4.46 divide each side by $0.12/mi
$0.12/0.12(M)=$4.46/$0.12/mi
M=37 mi charged miles in excess of 100 mi, so:
Total mileage=TM=Charged mileage(M)+100 mi
TM=M+100 mi=37 mi+100 mi=137 mi total mileage
ANSWER: Millie drove a total of 137 miles in the van.